OpenStudy (anonymous):

You are making poster board boundary markers in the shape of square pyramids for a relay race. If you are including a ten -inch square as the bottom and you want the slant height to be twelve inches, how much poster board do you need?

OpenStudy (bahrom7893):

what is a surface area of a square pyramid?

OpenStudy (anonymous):

\[SA= b ^{2} + 2bs\]

OpenStudy (bahrom7893):

SA=100+20*12=100+240=340

OpenStudy (anonymous):

how do u know what the base is

OpenStudy (anonymous):

bah did u leave